Tohban Report 2015-03-11

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 4 Mar 2015
End Date: 11 Mar 2015
Tohban: Lindsay Glesener
Tohban email:
Next Tohban: Juan Carlos Martínez Oliveros
List all reports


Solar Activity

FLARES!!! Awesome, occulted flares. And then loads of unocculted ones! There are only two ARs on the disk, but one is quite active. AR 12297 came out swinging, with decent-sized flares occurring before it even rotated onto the disk, and has been consistently producing ever since. A Major Flare Watch was called on March 7th in response to an M9.2 flare. There have been several M flares since, and the region has not quieted, so the Major Flare Watch is still on. Activity is expected to remain active at least for the near future. There was an X2 flare at ~16:20 UT today!

How many GOES flares occurred?

 Flares above B, C, M, X class were      8    45     8     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      6    30     7     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      0     0     0     0

There were RHESSI flares/GOES flares 322 / 61 over the time range 04-Mar-15 11-Mar-15

Memory Management

Decimation was changed from normal_vigorous to active vigorous at 065-19:17:19 UTC (March 6).

We turned off nighttime events starting 065-19:49:21 UTC (March 6).

Both of these actions were in response to a productive AR coming over the east limb in the south. This AR eventually produced a major flare watch (still ongoing as of March 11).

However, the memory is emptying, so we could turn nighttime events back on if desired.

Spacecraft Status

We spun up from 14.5 to 14.9 RPM between 2015-064-19:39 and 064-21:03 UTC (March 5).

Data Gaps

There were a number of gaps in SOH data:

 N_GAPS          13
 GAP START TIME              GAP END TIME                   GAP (SEC)
 2015-03-06T15:00:00.000 -- 2015-03-06T15:05:00.000       300.00000
 2015-03-06T15:20:00.000 -- 2015-03-06T15:25:00.000       300.00000
 2015-03-06T17:20:00.000 -- 2015-03-06T17:40:00.000       1200.0000
 2015-03-06T18:05:00.000 -- 2015-03-06T18:10:00.000       300.00000
 2015-03-06T18:30:00.000 -- 2015-03-06T18:35:00.000       300.00000
 2015-03-06T18:40:00.000 -- 2015-03-06T18:45:00.000       300.00000
 2015-03-06T18:50:00.000 -- 2015-03-06T18:55:00.000       300.00000
 2015-03-06T20:00:00.000 -- 2015-03-06T20:05:00.000       300.00000
 2015-03-06T20:40:00.000 -- 2015-03-06T20:45:00.000       300.00000
 2015-03-06T22:50:00.000 -- 2015-03-06T22:55:00.000       300.00000
 2015-03-06T23:10:00.000 -- 2015-03-06T23:15:00.000       300.00000
 2015-03-06T23:20:00.000 -- 2015-03-06T23:25:00.000       300.00000
 2015-03-06T23:45:00.000 -- 2015-03-07T00:00:00.000       900.00000
 GAP START TIME              GAP END TIME                   GAP (SEC)
 2015-03-07T00:25:00.000 -- 2015-03-07T00:30:00.000       300.00000


Detector issues

On March 5 we changed the Rear Fast threshold for detector nine to free up livetime eaten up by fast rates. Another change has been requested (and will happen today). Detector 9 seems to be deteriorating quickly in this respect.

On March 9 the D6 front fast threshold was changed due to low livetime:

Other notes

Spacecraft Management

Decimation Active/Vigorous
Night time data (fronts) plus/minus 4 minutes only
Night time data (rears) plus/minus 4 minutes only
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Normal
Personal tools